# Intrinsic Diophantine approximation: a solution to Mahler’s problem

### v1

@@ 5. Some perspectives
We expect that similar, but more refined, tools allow one to use arguments in the same spirit as in Section 3 in order to treat IFSs of the form
$T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m},$
where $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$. More precisely, we conjecture the following.
Conjecture 5.1.
Let $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m}$, where $q_{1},\dots,q_{m}\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, be a self-similar IFS, and let $K_{T}$ be its attractor. If $K_{T}$ has empty interior, then for any non-increasing function $\psi:\mathbb{N}\to\mathbb{R}_{+}$ one has
$\dim_{H}\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{for infinitely many }r\in\mathbb{Q}\cap K_{T}\Big\}=\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}},$
where $\delta_{\psi}=\liminf_{q\to+\infty}\frac{-\log\psi(q)}{\log q}$.
@@
with $q_{1},\dots,q_{m},t_{1},\dots,t_{m},p_{1},\dots,p_{m}\in\mathbb{Z}$ and $|t_{i}|<|q_{i}|$ for every $1\leq i\leq m$, we again suspect that a more refined argument in the same vein as in Section 3 may allow one to prove that, if $K_{T}$ has empty interior, then for any non-increasing $\psi:\mathbb{N}\to\mathbb{R}_{+}$, with $\delta_{\psi}$ as above,
$\dim_{H}\Big\{x\in K_{T}:|x-r|\leq\psi(h(r))\ \text{for infinitely many }r\in\mathbb{Q}\cap K_{T}\Big\}\leq\frac{\dim_{H}K_{T}}{\max\{1,\delta_{\psi}\}}.$
In this general case, the lower bound seems much more speculative. The reason is that, in general, rationals in such fractals no longer correspond to eventually periodic sequences. Thus, studying this problem requires a new description of the possible codings of rationals in this setting; see [23].
@@ 5.2. A more refined full-measure statement for homogeneous rational IFSs
Let $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q}+\frac{p_{i}}{q}\right\}$, where $q\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let $K_{T}$ denote its attractor. As mentioned in Section 2.2.2, these IFSs always satisfy the weak separation condition, and hence are Ahlfors-regular. By [22], this implies that
@@
$\begin{cases}\mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})=\mathcal{H}^{\dim_{H}K_{T}}(K_{T})\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}=+\infty,\\[4.30554pt] \mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})=0\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}<+\infty?\end{cases}$
Finally, we mention that another class of self-similar IFSs known to satisfy the WSC is given by Bernoulli convolutions associated with inverses of Pisot numbers. In this particular case, although the attractor is an interval, it is conjectured (see [6]) that all self-similar measures are non-Rajchman, which implies in particular that these measures are singular with respect to Lebesgue measure. Thus, in the above statement, one might in general be able to replace the Hausdorff measure $\mathcal{H}^{\dim_{H}K_{T}}$ by a self-similar measure of full dimension.
@@ 6. Appendix
6.1. Rational numbers in translates of one-dimensional homogeneous fractals
@@
Assume now that $\mathcal{L}(K)>0$. By the Besicovitch density theorem, there exist $x\in K$ and $r_{x}>0$ such that for any $0<r\leq r_{x}$,
@@
which implies

### v3

@@ 5.1. About the connection between equidistributuion mod q and Mahler’s problem
with $q_{1},\dots,q_{m},t_{1},\dots,t_{m},p_{1},\dots,p_{m}\in\mathbb{Z}$ and $1\leq|t_{i}|<|q_{i}|$ for every $1\leq i\leq m$. For such an IFS, it is easily checked any ultimately periodic sequence produces a rational. But, in these settings, it follows from [26] that it is no longer true in general that every rational in such fractals admits an eventually periodic coding. We refer to the Appendix, Subsubsection 6.1.2 for more details about this fact. Therefore, studying this problem requires a new description of the possible codings of rationals in this setting.
@@ 5.3. A more refined full-measure statement for homogeneous rational IFSs
Let $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q}+\frac{p_{i}}{q}\right\}$, where $q\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in\mathbb{Z}$, and let $K_{T}$ denote its attractor. As mentioned in Section 2.2.2, these IFSs always satisfy the weak separation condition, and hence are Ahlfors-regular (see Proposition 2.3). By [25], this implies that
@@
$\mathcal{H}^{\dim_{H}K_{T}}(E_{\psi,T,\text{int}})>0\iff\displaystyle\sum_{n\geq 1}nm^{n}\psi(q^{n})^{\dim_{H}K_{T}}=+\infty?$
Finally, we mention that another class of self-similar IFSs known to satisfy the WSC is given by Bernoulli convolutions associated with inverses of Pisot numbers. In this particular case, although the attractor is an interval, it is conjectured (see [7]) that all self-similar measures are non-Rajchman, which implies in particular that these measures are singular with respect to Lebesgue measure. Thus, in the above statement, one might not, in general, be able to replace the Hausdorff measure $\mathcal{H}^{\dim_{H}K_{T}}$ by a self-similar measure of full dimension.
@@ 6. Appendix
6.1. Rational numbers in self-similar sets
This sub-section aims at justifying that it is not generic in general for self-similar sets to contain rationals (this is the subject of the next subsubsection) and to make a short state of the art of the existence and possible codings of rational in self-similar sets associated with self-similar IFS defined by rational parameters.
6.1.1. Rational numbers in translations of one-dimensional homogeneous fractals
@@
Assume now that $\mathcal{L}(K)>0$. By Besicovitch density theorem [5, Theorem 3], there exist $x\in K$ and $r_{x}>0$ such that for any $0<r\leq r_{x}$,
@@
which implies, again by Besicovitch density theorem [5, Theorem 3], that
@@
6.1.2. Rationals and self-similar IFS’s defined by rational parameters
Let $\lambda_{1},...,\lambda_{m}\in\mathbb{Q}\cap\Big((-1,1)\setminus\left\{0\right\}\Big),$ $a_{1},....,a_{m}\in\mathbb{Q}$ be rationals and set
$T=\left\{f_{i}(\cdot)=\lambda_{i}\cdot+a_{i}\right\}_{1\leq i\leq m}.$
For every $n\in\mathbb{N},$ for every $0\leq k\leq n-1,$ for every $(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n},$
$\pi\Big((i_{1},...,i_{n}),(i_{k+1},...,i_{n})_{\infty}\Big)$
$=\sum_{1\leq t\leq n}\Big(\prod_{1\leq j<t}\lambda_{i_{j}}\Big)a_{i_{t}}+\Big(\prod_{1\leq j\leq n}\lambda_{i_{j}}\Big)\frac{\sum_{k+1\leq t\leq n}\Big(\prod_{k+1\leq j<t}\lambda_{i_{j}}\Big)a_{i_{j}}}{1-\prod_{k+1\leq j\leq n}\lambda_{i_{j}}}\in\mathbb{Q}.$
Hence $K_{T}$ always contains rational numbers. It is not true in general that, for any such $T$, rationals in $K_{T}$ corresponds to projections of ultimately periodic sequences. This is due to the following result.
Theorem 6.2.
Let $\lambda\in(0,1)$ and $T=\left\{0,...,\lfloor\frac{1}{\lambda}\rfloor\right\}.$ Then
$\mathbb{Q}(\lambda)\cap K_{T}\subset\left\{\pi\Big((i_{1},...,i_{n}),(i_{k+1},...,i_{n})_{\infty}\Big)\right\}_{0\leq k\leq n-1,n\in\mathbb{N},(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}}$
$\Longrightarrow(\frac{1}{\lambda}\text{ is Pisot or Salem}).$
In particular, for $\lambda=\frac{2}{3},$ $\mathbb{Q}(\lambda)=\mathbb{Q}$ and $\frac{1}{\lambda}$ is not an algebraic integer (thus is not Pisot nor Salem). Hence there must exist $r\in\mathbb{Q}\cap K_{T}$ which does not admit an ultimately periodic coding.
As mentioned in the introduction, in the special case where $T$ is of the form $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}_{1\leq i\leq m},$ where $q_{i},p_{i}$ are integers, it was proved by Fishman and Simmons , when $T$ satisfies the open set condition, that the rationals in $K_{T}$ corresponds to eventually periodic sequences. Later on, in [1], Baker remarked (but did not prove explicitly) that the assumption on the separation condition can be removed. For the sake of completeness, we prove this fact.
Proposition 6.3 ([19, 1]).
Let $q_{1},...,q_{m}\geq 2$ and $p_{1},...,p_{m}\in\mathbb{Z}$ be integers and set $T=\left\{f_{i}(\cdot)=\frac{\cdot}{q_{i}}+\frac{p_{i}}{q_{i}}\right\}.$ Then $r\in\mathbb{Q}\cap K_{T}$ if and only if there exists $n\in\mathbb{N},$ $0\leq k\leq n-1$ and $(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}$ such that
$x=\pi\Big((i_{1},...,i_{n},(i_{k+1},...,i_{n})_{\infty})\Big).$
Proof.
As $T$ has rational parameters, one only needs to prove that if $r\in\mathbb{Q}\cap K_{T},$ there exists $n\in\mathbb{N},$ $0\leq k\leq n-1$ and $(i_{1},...,i_{n})\in\left\{1,...,m\right\}^{n}$ such that
$r=\pi\Big((i_{1},...,i_{n},(i_{k+1},...,i_{n})_{\infty})\Big).$
Fix $r=\frac{a}{b}\in\mathbb{Q}\cap K_{T}$ and $(j_{n})_{n\in\mathbb{N}}\in\left\{1,...,m\right\}^{\mathbb{N}}$ such that
$\pi\Big((j_{n})_{n\in\mathbb{N}}\Big)=r.$
For any $1\leq i\leq m,$
$f_{i}^{-1}(r)=q_{i}r-p_{i}\in\frac{\mathbb{Z}}{b}.$
Hence, for any $t\in\mathbb{N},$ $f_{j_{t}}^{-1}\circ...\circ f_{j_{1}}^{-1}(r)\in K_{T}\cap\frac{\mathbb{Z}}{b}.$ Thus there exists $n\in\mathbb{N}$ and $0\leq k\leq n-1$ such that
$f_{j_{n}}^{-1}\circ...f_{j_{i}}^{-1}(r)=f_{j_{k}}^{-1}\circ...f_{j_{i}}^{-1}(r).$
In particular, $x=f_{j_{k}}^{-1}\circ...\circ f_{j_{1}}^{-1}(r)$ is a fixed point of $f_{j_{n}}^{-1}\circ...\circ f_{j_{k+1}}^{-1},$ hence is a fixed point of $f_{j_{k+1}}\circ...\circ f_{j_{n}}.$ Writing $g=f_{j_{k+1}}\circ...\circ f_{j_{n}},$
$\lim_{t\to+\infty}g^{(t)}(0)=x,$
where $g^{(t)}$ is $g$ composed $t$ times. This yields
$x=\pi\Big((j_{k+1},...,j_{n})_{\infty}\Big)\Rightarrow r=\pi\Big((j_{1},...,j_{k},(j_{k+1},...,j_{n})_{\infty})\Big),$
which proves the claim. ∎
@@ 6.2. WSC and Ahlfors regularity
6.3. About Remark 1.2
In this section, we prove the following.
Proposition 6.4.
Let $T$ be a self-similar IFS of the form
$T=\Big\{f_{i}(\cdot)=\frac{\cdot}{b}+\frac{p_{i}}{b}\Big\}_{1\leq i\leq m},$
where $b\geq 2$ and $p_{1},....,p_{m}\in\mathbb{Z}$ and write $K_{T}$ its attractor. The three following items are equivalent:
  • (1)
    $\mathcal{L}(K_{T})>0,$
  • (2)
    $K_{T}$ has non empty interior,
  • (3)
    $\dim_{H}K_{T}=1.$
Proof.
Call again $\widetilde{K}_{T}\subset\mathbb{T}^{1}$ the projection of $K_{T}$ on $\mathbb{T}^{1}$ and recall that $\widetilde{K}_{T}$ is $\times b$ invariant. Clearly $(2)\Rightarrow(1).$
We prove $(1)\Rightarrow(2):$ Assume that $\mathcal{L}(K_{T})>0$ and notice that $\mathcal{L}(K_{T})>0\Rightarrow\mathcal{L}(\widetilde{K}_{T})>0.$ Since $\mathcal{L}$ is $\times b$ ergodic, there exists $x\in\mathbb{T}^{1}$ such that
$\lim_{n\to+\infty}\frac{\sum_{k=0}^{n-1}\chi_{\widetilde{K}_{T}}(b^{k}x)}{n}=\mathcal{L}(\widetilde{K}_{T}),$
where $\chi_{\widetilde{K}_{T}}$ is the indicator function of $\widetilde{K}_{T}$. As $y\in\widetilde{K}_{T}\Rightarrow by\in\widetilde{K}_{T},$ one has
$\lim_{n\to+\infty}\frac{\sum_{k=0}^{n-1}\chi_{\widetilde{K}_{T}}(b^{k}x)}{n}=1$
so that $\mathcal{L}(\widetilde{K}_{T})=1$. $\widetilde{K}_{T}$ being compact, this yields $\widetilde{K}_{T}=\mathbb{T}^{1}$ and $K_{T}$ has non empty interior.
As $(1)\Rightarrow(3)$ and by Proposition 2.3, $K_{T}$ is $\dim_{H}K_{T}$-Alfhors regular, $(1)\Leftrightarrow(3).$
Acknowledgments: The author would like to thank J. Barral and S. Seuret for their valuable advice regarding this manuscript, as well as B. Wang for pointing out a mistake in the `` proof" of Conjecture 1.4.